A simple pendulum has a length of 1.5m and a bob of mass 0.2kg. If the amplitude of oscillation of the bob is 0.05m, calculate period of oscillation
1
Expert's answer
2020-01-31T10:39:24-0500

The period of oscillation:


"T=2\pi \sqrt{\frac{l}{g}}"

"T=2\pi \sqrt{\frac{1.5}{9.8}}=2.5\ s"